OpenStudy (anonymous):

I have another one to solve U-3= 1/125 U=

pooja195 (pooja195):

add 3 to 1/125 so 1/125+3=?

pooja195 (pooja195):

HINT : your answer will have a unit

pooja195 (pooja195):

hello?

pooja195 (pooja195):

ok well..... \(\huge\color{black}{WELCOME~TO~OPENSTUDY!!!~:)}\)

OpenStudy (anonymous):

1/128

pooja195 (pooja195):

no we want units so the fraction will stay the same your just adding 3

pooja195 (pooja195):

dont think too hard ;)

OpenStudy (anonymous):

so then do you -3-125?

pooja195 (pooja195):

nope just ADD 3

pooja195 (pooja195):

i told you this 1/125 will remain the same =)

OpenStudy (anonymous):

u=128

pooja195 (pooja195):

nope remember UNITS

OpenStudy (anonymous):

1/125 = +3

pooja195 (pooja195):

nope add three so x 1/125 what is X is your adding 3?

OpenStudy (anonymous):

375/125

OpenStudy (anonymous):

4/125

pooja195 (pooja195):

nope 3 1/125 because we're adding 3 thats it nothing needs to be done

OpenStudy (dan815):

a number - 3 is 1/125 so so this number must be 3+1/125

OpenStudy (anonymous):

so then it would be u= 3+1/125

OpenStudy (dan815):

right, now simplify it

OpenStudy (anonymous):

how do I do that?

OpenStudy (dan815):

represent it as just 1 term, you can add the numerators if they have the same denominator

OpenStudy (dan815):